# CGD: Modifying the Loss Landscape by Gradient Regularization

### v2

@@ 5 Reinterpretation of Explicit Gradient Regularization (EGR)
EGR has the drawbacks of having fictitious minima identified in Lemma 1. Since, the loss landscape of neural networks is highly uneven [11], it’s likely that artificial stationary points are introduced and the optimizer might converge at local-maxima points since their nature changes with regularization. For example, in Fig. 5 (Right) we observe artificial stationary points being introduced between local-minima and local-maxima of the original loss function while some local-maxima also turn into local-minima of the penalized loss function. Therefore, one needs to ensure that there are suitable fixes for these scenarios for better performance. Specifically, the direction along of the penalized loss function, $-\nabla g(\cdot)$ should be a descent direction and that we shouldn’t stop at points where $\nabla g(\cdot)={\mathbf{0}}$ but $\nabla f(\cdot)\not={\mathbf{0}}$ as discussed in Lemma 1.
@@
6 Future Works
In this work, we considered a new line search method that penalizes the norm of the gradient and provides iterates that have lower gradient norm compared to vanilla gradient descent. We identify properties of this algorithm, provide a variant that does not require Hessian and illustrate connections to the widely popular explicit gradient regularization literature.
There are several future directions arising from this work. We would like to investigate in greater detail the role of different penalty functions $h(\cdot)$. We would also like to investigate applications of this method in more diverse settings like reinforcement learning and even Bayesian optimization. We hope this work sparks more discussions on gradient regularization helping in neural network generalization.
@@ 6.0.1 Acknowledgements
Appendix 0.A How CGD affects the Condition Number of Quadratic Function
Let the quadratic function be $f\left(\mathbf{x}\right)=\frac{1}{2}\mathbf{x}^{T}Q\mathbf{x}-\mathbf{b}^{T}\mathbf{x}$ where $Q\succ 0$. Let $l\triangleq\lambda_{\min}(Q)$ and $L\triangleq\lambda_{\max}(Q)$.
From Equation 9 we know $\mathbf{x}_{k+1}=\mathbf{x}_{k}-\alpha B_{k}\nabla f_{k}$ where $B_{k}={I}+2\lambda Q$ and $\nabla f\left(\mathbf{x}_{k}\right)=Q\mathbf{x}_{k}-\mathbf{b}$. Thus, the error at each iterate can be given as:
[…]
$\leq\left\|{{I}-\alpha B_{k}Q}\right\|\left\|\mathbf{e}_{k}\right\|$ (14)
Since ${I}$ and $Q$ commute, we have $(1+2\lambda l){I}\preccurlyeq B_{k}\preccurlyeq(1+2\lambda L){I}$. Also since both $Q$ and $B_{k}$ are symmetric positive-definite matrices, $\|B_{k}Q\|$ has the following bounds:
$\lambda_{\min}\left(B_{k}\right)\lambda_{\min}(Q)\leq\left\|B_{k}Q\right\|\leq\lambda_{\max}\left(B_{k}\right)\lambda_{\max}(Q)$ (15)
$\therefore\left\|{{I}-\alpha B_{k}Q}\right\|\leq\left(1-\alpha\lambda_{\min}\left(B_{k}\right)\lambda_{\min}(Q)\right)=1-\alpha(1+2\lambda l)l$
So in Equation 14,
$\left\|{\mathbf{x}_{k+1}-\mathbf{x}^{\ast}}\right\|$ $\leq\left(1-\alpha(1+2\lambda l)l\right)\left\|{\mathbf{x}_{k}-\mathbf{x}^{\ast}}\right\|$
$\leq\left(1-\alpha(1+2\lambda l)l\right)^{k}\left\|{\mathbf{x}_{0}-\mathbf{x}^{\ast}}\right\|$
From Equation 15, substitute $\alpha$ as:
$\Rightarrow\left\|{\mathbf{x}_{k+1}-\mathbf{x}^{\ast}}\right\|$ $\leq\left(\frac{(1+2\lambda L)L-(1+2\lambda l)l}{(1+2\lambda L)L+(1+2\lambda l)l}\right)^{k}\left\|{\mathbf{x}_{0}-\mathbf{x}^{\ast}}\right\|$
where ${\kappa}=\kappa(B_{k})\kappa(Q)$ where $\kappa(A)\triangleq\frac{\lambda_{\max}{(A)}}{\lambda_{\min}(A)}$ is the condition number of matrix $A$.

### v3

@@ 5 Reinterpretation of Explicit Gradient Regularization (EGR)
Two curves show the original loss landscape and the penalized landscape. Overlaid on these are two optimization trajectories for GD and CGD, respectively. The CGD path explores a broader region and reaches a lower minimum than the GD path, illustrating improved descent behavior by following directions computed from the modified (penalized) loss function.
Two curves showing the original loss landscape and the penalized landscape. The penalized curve slightly deviates from the original, with several regions where local maxima in the loss landscape correspond to local minima in the penalized version.
@@
EGR has the drawbacks of having fictitious minima identified in Lemma 1. Since, the loss landscape of neural networks is highly uneven [12], it’s likely that artificial stationary points are introduced and the optimizer might converge at local-maxima points since their nature changes with regularization. For example, in Fig. 5 (Right) we observe artificial stationary points being introduced between local-minima and local-maxima of the original loss function while some local-maxima also turn into local-minima of the penalized loss function. Therefore, one needs to ensure that there are suitable fixes for these scenarios for better performance. Specifically, the direction along of the penalized loss function, $\nabla g(\cdot)$ should be a descent direction and that we shouldn’t stop at points where $\nabla g(\cdot)={\mathbf{0}}$ but $\nabla f(\cdot)\not={\mathbf{0}}$ as discussed in Lemma 1.
@@
6 Conclusions and Future Work
In this work, we considered a new line search method that penalizes the norm of the gradient and provides iterates that have lower gradient norm compared to vanilla gradient descent. We identify properties of this algorithm, provide variants that do not require the Hessian and illustrate connections to the widely popular explicit gradient regularization literature.
There are several future directions arising from this work. We would like to investigate the role of different penalty functions $h(\cdot)$ in greater detail. Another future direction is using a combination of negative and positive $\{\lambda_{k}\}$ values in our iterates. Particularly, negative $\{\lambda_{k}\}$ values might help in descending faster at points where the Hessian is negative definite. We would also like to explore applications of our method in more diverse settings like reinforcement learning (RL) and Bayesian optimization (BO). In particular, for RL, we can develop penalized variants of existing Policy Gradient Methods [21, 20, 18], which can be evaluated to assess the effectiveness of gradient regularization in RL settings. Similarly, for BO tasks, we can construct acquisition functions involving a gradient penalty, in a manner similar to [16], which might improve performance. We hope this work sparks further discussion on gradient regularization helping in neural network generalization.